OpenStudy (anonymous):

Yes. See the value of y when x = 8 or x = -8. Same with other values. You need to calculate based on the equation and see whether it matches.

OpenStudy (anonymous):

I have one last question x - 2= 1/8(y+1)^2 how will that look like in a graph? because i try graphing in a graph er and it says error ..is there a way i can? :S

OpenStudy (anonymous):

You need to express it as: y = something.

OpenStudy (anonymous):

x - 2= 1/8(y+1)^2 is this the graph http://i729.photobucket.com/albums/ww292/akatushi101/ayooamama.jpg